OpenStudy (anonymous):

An 800 g rock is tied to a string and whirled in a vertical circle with a 75 cm radius and a frequency of rotation of 1.6 Hz. a. Assuming a constant velocity, determine the tension in the rope at the top of the circle. b. Assuming a constant velocity, determine the tension in the rope at the bottom of the circle.

OpenStudy (anonymous):

you still need help?

OpenStudy (anonymous):

yes please!

OpenStudy (anonymous):

i'll help you but just spare me a minute ...

OpenStudy (anonymous):

Guess.. @PFEH.1999 must be moving close to speed of light.. cause his 1 minute is still not up and our 10 hours are up lolzz JK

OpenStudy (anonymous):

can u find the required centripetal force or acceleration ?

OpenStudy (anonymous):

OH yes @Mashy , i forgot all about this one :D i went to solve a question in math and i forgot ... sry :) :D

OpenStudy (anonymous):

would you be able to help me now @PFEH.1999 please?

OpenStudy (anonymous):

T-0.8kg*9.81m/s2=0/8kg*4pi2*1.6Hz2*0.75m T=0.8kg[(4pi2*1.6Hz2*0.75m)-9.81m/s2] = 52.8N or at least tell me whether or not that is correct?